the perimeter of a rectangle is 110 inches the length is 10m more than twice the width find the dimensions
Answer by checkley77(12844)

P=2L+2W
L=2W+10
REPLACE L WITH (2W+10) IN THE PERIMEYTER EQUATION & SOLVE FOR W
110=2(2W+10)+2W
110=4W+20+2W
110=6W+20
6W=110-20
6W=90
W=90/6
W=15 ANS. FOR THE WIDTH.
L=2*15+10
L=30+10=40 ANS FOR THE LENGTH.
PROOF:
110=2*40+2*15
110=80+30
110=110